# FreezeAsGuard: Mitigating Illegal Adaptation of Diffusion Models via Selective Tensor Freezing

### v1

@@ 4.1 Quantitative Results of Mitigation Power in Illegal and Innocent Domains
GPU Memory (GB) 18.28 18.26 16.97 16.96 15.43 14.15 13.61 9.49
Per-batch computing time (s) 1.17 1.14 1.09 1.06 1.05 1.02 1.00 0.91
@@
At the same time, as shown in Table 3, by applying FreezeAsGuard’s selection of tensor freezing, users can save 22%-48% GPU memory and 13%-21% wall-clock computing time, compared to other baselines without any freezing ($\rho$=0%). Such savings, hence, well motivate users to adopt the FreezeAsGuard’s tensor freezing in their fine-tuning practices.
When the freezing ratio ($\rho$) increases, the difference between FreezeAsGuard and random freezing diminishes, and their mitigation powers also reach the same level. This means that only a small set of tensors are important for adaptation in a specific domain. With a high freezing ratio, random freezing is more likely to freeze these important tensors. However, at the same time, it could also freeze tensors that are important to innocent domains, resulting in low adaptation performance in innocent domains.
Based on these results, we empirically conclude that $\rho$=30is the optimal freezing ratio on SD v1.5, to ensure sufficient mitigation power in illegal domains and the minimum impact in innocent domains. For different illegal domain scales (i.e., different numbers of subjects involved) and SD models, we will later demonstrate that the optimal freezing ratio generally stays between 20% and 30%.
Figure 7: Qualitative examples of generated images in illegal domains of 10 subjects in FF25 dataset, after applying FreezeAsGuard-30% to fine-tuning SD v1.5. Each prompt adopts the same seed for generation. More image examples are in Appendix G.
4.2 Qualitative Examples of Generated Images
Most existing metrics cannot precisely reflect the image quality perceived by humans [31; 17]. Instead, being similar to those shown in Figure 2, we show more qualitative image examples in Figure 7, which show that FreezeAsGuard effectively prevents the generated images from being recognized as the target subjects by human perception. Meanwhile, the fine-tuned model can still generate detailed background content and subjects’ postures aligned with the prompt, indicating that the mitigation power is highly selective and focuses only on subjects’ faces. In contrast, although UCE [25] and IMMA [59] can unlearn the knowledge of illegal domains by modifying the model weights, the illegal user can still relearn such knowledge via fine-tuning and the fine-tuned model can still generate images with decent quality. The key reason is that these unlearning methods do not have any restriction on the fine-tuning process, which instead, can be effectively regulated by FreezeAsGuard via tensor freezing.
Besides, freezing more model tensors could reduce overfitting in illegal domains and possibly improve the quality of fine-tuning. As shown in Figure 8, we pick some examples that can be easily mitigated by FreezeAsGuard with a very low $\rho$ (e.g., 1%). Then, when $\rho$ increases from 1% to 80%, FreezeAsGuard exhibits constant mitigation power despite the potentially less model overfitting. This is because the overfitting reduction is implicitly incorporated into mask learning by simulating different selections of tensor freezing, and the mitigation power hence overweighs the overfitting reduction.
Figure 8: Qualitative image examples after applying FreezeAsGuard to fine-tuning SD v1.5, with different freezing ratios in illegal domains. Each row adopts the same seed for generation.
4.3 Mitigation Power with Different Illegal Domain Scales
We randomly pick images from 2, 5, and 10 subjects in FF25 dataset, as the illegal domain. As shown in Table 4, FreezeAsGuard can reduce the FID of generated images by up to 14% compared to full FT. In comparison, UCE exhibits no mitigation power in some cases since fine-tuning can relearn the unlearned knowledge in illegal domains. IMMA aggressively modifies the weights and results in significant performance drops in innocent domains. With more subjects in illegal domain, the difference of mitigation between FreezeAsGuard and random freezing becomes smaller, because more subjects correspond to more adaptation-critical tensors, and random freezing is more likely to cover them.
4.4 Different Innocent Domains
It is possible that illegal domains and innocent domains share some critical parameters in common. However, as shown in Table 5, FreezeAsGuard maximally retains the adaptation performance in different innocent domains, with on-par or even better image quality compared to UCE and IMMA. In some cases, the CLIP scores of generated images are better than the training samples, possibly because the fine-tuned model finds a shortcut that deceives the pre-trained CLIP model. We are then motivated to check the qualitative results as shown in Figure 9. The generated images are either similar to training samples, or more colorful and contain more details. Besides, the generated images are not perfectly aligned with prompts, and we attribute these errors to the limitation of open-sourced SD models. These errors can be mitigated with other larger diffusion models, such as DALL-E 3 [10] and Imagen2 [12].
Method 2 subjects (seed 1) 2 subjects (seed 2) 5 subjects 10 subjects
Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$) Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$) Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$) Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$)
Full FT 139.20 32.83 135.61 32.83 127.38 32.83 143.58 32.83
UCE 138.39 30.74 147.35 30.27 131.27 30.27 151.53 30.25
IMMA 139.88 30.03 145.39 29.81 129.60 28.99 148.05 28.71
Random-20% 139.04 29.80 142.72 29.80 135.48 29.80 152.65 29.80
FreezeAsGuard-20% 151.30 30.31 153.90 30.10 135.85 31.12 156.09 30.26
Random-30% 138.16 28.92 142.64 28.92 136.65 28.92 152.21 28.92
FreezeAsGuard-30% 146.30 31.49 153.00 30.29 138.09 30.48 155.98 30.47
Table 4: Mitigation power in illegal domains with different scales (as numbers of subjects in FF25 dataset) and innocent domains (Logo dataset)
Method Logo dataset Clothes dataset
CLIP ($\uparrow$) TOPIQ ($\uparrow$) CLIP ($\uparrow$) TOPIQ ($\uparrow$)
GT 31.24 - 29.63 -
Full FT 32.83 0.097 32.84 0.059
UCE 30.27 0.088 32.47 0.070
IMMA 28.99 0.097 31.01 0.065
R-10% 30.28 0.091 32.33 0.055
FG-10% 31.41 0.100 31.98 0.065
R-20% 30.36 0.093 32.44 0.064
FG-20% 31.12 0.103 31.73 0.062
Table 5: Results on different innocent data domains with photos of 5 subjects as the illegal domain (GT: Ground Truth, FG: FreezeAsGuard)
Figure 9: Generated image examples on the logo dataset as the innocent domain
4.5 Different Diffusion Models
We apply FreezeAsGuard to three different SD models. As shown in Table 6, FreezeAsGuard constantly outperforms baseline schemes on all diffusion models. According to the full fine-tuning performance, SD v1.4 and v1.5 are generally stronger than v2.1. Accordingly, FreezeAsGuard’s retaining power in the innocent domain is slightly better for v1.4 and v1.5 models. We hypothesize that better pre-trained models have more modularized knowledge distribution over model parameters, and hence allow FreezeAsGuard to freeze them without affecting innocent domains.
Method Stable Diffusion v1.4 Stable Diffusion v1.5 Stable Diffusion v2.1
Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$) Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$) Illegal FID ($\downarrow$) Innocent CLIP ($\uparrow$)
Full FT 146.03 33.26 143.58 32.83 148.37 32.44
UCE 151.06 30.12 151.53 30.27 152.77 27.48
IMMA 150.49 29.07 148.05 28.71 149.49 26.89
Random-20% 151.42 30.00 152.65 29.80 152.52 28.43
FreezeAsGuard-20% 156.11 30.27 156.09 30.26 155.47 28.47
Table 6: Quantitative results on illegal (10 subjects in FF25 dataset) and innocent domains (Logo dataset) using different diffusion models
@@ 5 Conclusion & Broader Impact
In this paper, we present FreezeAsGuard, a new technique for mitigating illegal adaptation of diffusion models by freezing model tensors that are adaptation-critical only for illegal domains. FreezeAsGuard largely outperforms existing model unlearning schemes. Our rationale for tensor freezing is generic and can be applied to other large generative models.
@@ Appendix A Vectorizing the Gradient Calculations in Bilevel Optimization
$\frac{\partial{\mathcal{L}_{upper}}}{\partial{w_{i}}}=\left<\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\frac{\partial{\theta(m)_{i}}}{m_{i}}\right>\frac{\partial m_{i}}{\partial w_{i}}=\left<\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\theta_{pre}^{(i)}-\theta_{ft}^{(i)}\right>\frac{1}{T}\sigma\left(\frac{w_{i}}{T}\right)\sigma\left(1-\frac{w_{i}}{T}\right),$ (11)
@@ Appendix B Deciding the Number of Fine-tuning Iterations in Bilevel Optimization
Appendix C Details of the Famous-Figures-25 (FF25) Dataset
Our FF25 dataset contains 8,703 portrait images of 25 public figures and the corresponding text descriptions. All the images were crawled from publicly available sources on the Web. These 25 subjects include politicians, movie stars, writers, athletes and businessmen, with diverse genders, races, and career domains. As shown in Figure 11, the dataset contains 400-1,300 images of each subject.
@@ Appendix C Details of the Famous-Figures-25 (FF25) Dataset
When crawling images on the web, we only consider images that 1) has a resolution higher than 512$\times$512 and 2) contains $>$3 faces detected by OpenCV face recognition API [13] as valid. Each raw image is then center-cropped to a resolution of 512$\times$512. For each image, we use a pre-trained BLIP2 image captioning model [38] to generate the corresponding text description, and prompt BLIP2 with the input of “a photo of <person_name> which shows” to avoid hallucination. For example, “a photo of Cristiano Ronaldo which shows”, when being provided to the BLIP2 model as input, could result in text description of “a photo of Cristiano Ronaldo which shows him smiling in a hotel hallway”. More sample images and their corresponding text descriptions are shown in Figure 12.
Appendix D Source Code Repository
We have packed all the source codes of FreezeAsGuard into the a .zip file and provided the file in the supplementary material. Please refer to the README.md file in the code repository for details about the source codes. In the README.md file, we also provided an anonymous link to our self-collected FF25 dataset.
Appendix E Sample Images of Datasets as Innocent Domains
The Logo dataset [4] contains minimalist designs of logo images with highly detailed text descriptions. As shown in Figure 14, the logos can correspond to different types of businesses, such as restaurants, coffee shops, and e-shops. The text description specifies the shape, color, and position of every object that should be included in the logo. Most images are near square and the logo is centered, and these images can hence be safely center-cropped to the standard resolution of 512$\times$512 for fine-tuning (being similar to those in the FF25 dataset).
Similarly, the Clothes dataset [3] contains images of fashion products with highly detailed text descriptions. As shown in Figure 14, the products include clothes (e.g., coats and T-shirts) and other wearings (e.g., scarves). The text description can specify not only direct visual information (e.g., color and shape) but also texture information (e.g., soft cotton and recycled wool). Despite having diverse resolutions, most products are centered on the images and can be first center-cropped and then rescaled to achieve 512$\times$512 for fine-tuning (being similar to those in the FF25 dataset).
Figure 13: Examples in Logo dataset
Figure 14: Examples in Clothes dataset
Appendix F Details of Evaluation Setup
@@
For subjects in illegal domains (i.e., public figures in the FF25 dataset) and innocent domains (i.e., subjects in Logo and Clothes datasets), we use 100 images of each subject for mask learning. The remaining data samples in the FF25 dataset are used for illegal model fine-tuning. Note that, to mitigate model adaptation in a specific illegal domain (e.g., a specific public figure), we will need to use data samples in the same domain for mask learning. However, in our evaluations, the set of data samples used for mask learning and the set of data samples used for illegal model fine-tuning never have any overlap. For example, to mitigate the fine-tuned model’s capability of generating portrait images of Barack Obama, we will use a set of portrait images of Barack Obama to learn the mask for tensor freezing. Then, another set of Barack Obama’s portrait images are used to emulate illegal users’ fine-tuning the diffusion model, and FreezeAsGuard’s performance of mitigating illegal model adaptation is then evaluated by the quality of images generated by the fine-tuned model regarding this subject.
@@
For mask learning, we set the gradient step size to 10, the simulated user learning rate to 1e-5, and iterate sufficient steps with the batch size of 16. The temperature for the mask’s continuous form is set to 0.2, which we empirically find to ensure sufficient sharpness without impairing trainability. When fine-tuning the diffusion model as an illegal user, we adopt a learning rate of 1e-5 and the batch size of 4 with Adam [35] optimizer, to fine-tune 2,000 iterations on subject’s data samples. Following the standard sampling setting of diffusion models, the loss is only calculated from a random denoising step during fine-tuning for every iteration, to ensure training efficiency. For image generation, we adopt the PNDMScheduler [34] and proceed with 50 denoising steps for sufficient image quality.
Figure 15: Qualitative image generation performance after applying FreezeAsGuard-30% to Stable Diffusion v1.5 on illegal domains, where each prompt adopts the same seed for generation
@@
Appendix G More Qualitative Examples of Images Generated by the Fine-tuned Model
@@
The image quality degradation introduced by FreezeAsGuard can be from different perspectives. As shown in Figure 15, in most cases, such as for Angela Merkel and Barack Obama, the distortion is exhibited as stretched faces or exaggerated emotions which make the subject unrecognizable. In some other cases, such as the third row of Cristiano Ronaldo’s photos, the generated image contains unrealistic duplication of subjects. Moreover, for the third row of Bill Gates’s photos, the subject in the generated image with FreezeAsGuard is not wearing a mask, which is not aligned with the prompt. This is because, with FreezeAsGuard’s tensor freezing, the model cannot correctly convert the text features extracted by the text encoder to the aligned image tokens.
Appendix H The Learned Selection of Frozen Tensors
In Figure 16, we present the learned masks with different freezing ratios on 10 subject’s portrait images as the illegal domain and the logo dataset as the innocent domain. The distribution of the frozen tensors is never uniform over all freezing ratios. We found that most of the selected tensors are the kernel tensors of conv_in layers in the UNet. As the freezing ratio increases, more bias tensors are selected to freeze.
(a) freezing ratio = 1%
(b) freezing ratio = 5%
(c) freezing ratio = 10%
(d) freezing ratio = 20%
(e) freezing ratio = 30%
(f) freezing ratio = 40%
Figure 16: Learned binary masks with different freezing ratios (1 means to freeze)

### v2

@@ 4.7 Reduction of Computing Costs
GPU Memory (GB) 18.28 18.26 16.97 16.96
Per-batch computing time (s) 1.17 1.14 1.09 1.06
[…]
GPU Memory (GB) 15.43 14.15 13.61 9.49
Per-batch computing time (s) 1.05 1.02 1.00 0.91
@@ 5 Conclusion & Broader Impact
In this paper, we present FreezeAsGuard, a new technique for mitigating illegal adaptation of diffusion models by freezing model tensors that are adaptation-critical only for illegal classes. FreezeAsGuard largely outperforms existing model unlearning schemes. Our rationale for tensor freezing is generic and can be applied to other large generative models.
Figure 12: The frozen tensors for illegal classes on the FF-25 dataset, with $\rho$=30%
Figure 13: The frozen tensors for illegal classes on the Artwork dataset, with $\rho$=70%
@@ Appendix A Vectorizing the Gradient Calculations in Bilevel Optimization
$\frac{\partial{\mathcal{L}_{upper}}}{\partial{w_{i}}}$ =$\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\frac{\partial{\theta(m)_{i}}}{m_{i}}\right>\frac{\partial m_{i}}{\partial w_{i}}$ (11)
=$\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\theta_{pre}^{(i)}-\theta_{ft}^{(i)}\right>\frac{1}{T}\sigma\left(\frac{w_{i}}{T}\right)\sigma\left(1-\frac{w_{i}}{T}\right),$ (12)
@@ Appendix B Deciding the Number of Fine-tuning Iterations in Bilevel Optimization
Appendix C Details of Datasets
The Famous-Figures-25 (FF25) Dataset: Our FF25 dataset contains 8,703 portrait images of 25 public figures and the corresponding text descriptions. These 25 subjects include politicians, movie stars, writers, athletes and businessmen, with diverse genders, races, and career domains. As shown in Figure 15, the dataset contains 400-1,300 images of each subject.
All the images were crawled from publicly available sources on the Web, using the AutoCrawler tool [4]. We only consider images that 1) has a resolution higher than 512$\times$512 and 2) contains $>$3 faces detected by OpenCV face recognition API [12] as valid. Each raw image is then center-cropped to a resolution of 512$\times$512. For each image, we use a pre-trained BLIP2 image captioning model [39] to generate the corresponding text description, and prompt BLIP2 with the input of “a photo of <person_name> which shows” to avoid hallucination. For example, “a photo of Cristiano Ronaldo which shows”, when being provided to the BLIP2 model as input, could result in text description of “a photo of Cristiano Ronaldo which shows him smiling in a hotel hallway”. We empirically find that adopting this input structure to the BLIP2 model produces much fewer irrelevant captions. More sample images and their corresponding text descriptions are shown in Figure 16.
@@ Appendix C Details of Datasets
The Artwork Dataset: We selected five renowned digital artists, each of which has a unique art style, and manually downloaded 100–300 representative images from their Instagram accounts. The total amount of images in the dataset is hence 1,134. We then used a pre-trained BLIP2 image captioning model [39] to generate text prompts for each image. In Figure 17, we show a sample image and its text prompt for each artist.
Figure 17: Examples of collected painting from 5 artists
The NSFW-Caption Dataset: This dataset contains 2,000 NSFW images collected from MetArt, and each image has a very detailed caption, as shown in Figure 18.
Figure 18: One sample in the NSFW-Caption dataset
Also, in evaluations of FreezeAsGuard’s capability of mitigating the generation of explicit contents, we use the Modern-Logo-v4 dataset [5], which contains 803 logo images that are labeled with informative text descriptions, as the legal class. As the examples in Figure 19 shown, these logos are minimalist, meeting modern design requirements and reflecting the corresponding company’s industry.
Figure 19: Examples in the Modern-Logo-v4 dataset
Appendix D Details of Image Quality Metrics
D.1 Domain-specific feature extractor
In general, we measure the quality of images generated by the fine-tuned diffusion model by comparing their similarity with the original training images used to fine-tune the diffusion model. Most commonly used image similarity metrics, such as FID [28], LPIPS [31] and CLIP score [27], compute the similarity between the distributions of the extracted features from the generated and original images [43, 27]. The feature vectors are obtained using image feature extractors like the Inception model [28]. They often perform reasonably well in measuring similarity between images of common objects, such as those included in the ImageNet data samples [48].
However, existing studies find that these metrics cannot reliably measure the similarity between very similar subjects, such as human faces of different human subjects or artworks in different art styles [30, 54]. In practice, we observe that the measured image quality by these metrics could even contradict human perception. For example, as shown in Figure 20, while images generated with FreezeAsGuard are significantly lower in quality and differ more from the training images from a human perspective, the LPIPS scores of images generated by the fully fine-tuned model (without applying FreezeAsGuard) are similar to ours, even though they look quite different visually.
Figure 20: Evaluating the similarity in art style using the LPIPS score [31], where a higher score means more difference from the original training image.
Therefore, to address the limitations of these generic image quality metrics, as described in the paper, we use domain-specific feature extractors to obtain features from the training and generated images, then compute the cosine distance between the feature vectors as the final measure of the generated images’ quality. For human faces, we select three top feature extractors, namely FaceNet-512 (FN-L), FaceNet (FN), and VGG-Face (VGG), as provided in the DeepFace package [50]. For art styles in artworks, we use a pretrained CSD model from [52].
D.2 NudeNet score
We use a NSFW detector, namely NudeNet [2], to decide if the generated images contain any explicit content. For an input image, NudeNet can output a list of detected human body parts (such as ANUS_EXPOSED and FACE_FEMALE), along with the corresponding probabilities of these body parts’ appearances in the image. We sum all these probabilities together as the NudeNet score of the image, with a lower score indicating a lower probability of containing explicit content. The full list of the detectable human body parts is as follows:
FEMALE_GENITALIA_COVERED,FACE_FEMALE,
BUTTOCKS_EXPOSED,FEMALE_BREAST_EXPOSED,
FEMALE_GENITALIA_EXPOSED,
MALE_BREAST_EXPOSED,ANUS_EXPOSED,
FEET_EXPOSED,BELLY_COVERED,FEET_COVERED,
ARMPITS_COVERED,ARMPITS_EXPOSED,FACE_MALE,
BELLY_EXPOSED,MALE_GENITALIA_EXPOSED,
ANUS_COVERED,FEMALE_BREAST_COVERED,
BUTTOCKS_COVERED,
and we select the following 5 from them as indicators of explicit content:
UTTOCKS_EXPOSED,FEMALE_BREAST_EXPOSED,
FEMALE_GENITALIA_EXPOSED,ANUS_EXPOSED,
MALE_GENITALIA_EXPOSED
D.3 Details of Human Evaluations
Our human evaluation involves 16 participants of college students. These participants ranged in age from 19 to 28, with 14 identifying as male and 2 as female. We conduct our human evaluation by distributing the images being examined by participants via an online questionnaire, which consists of multiple sets of images. In each set of images, a training image is first shown as a reference, and then several images generated by the fine-tuned diffusion models in different ways (e.g., unprotected full fine-tuning, UCE, IMMA, FreezeAsGuard) are shown, with respect to the same text prompt. The participants are asked to rate each generated image based on how closely it resembles the same subject (public figures or art styles) as shown in the reference image. The rating scale ranges from 1 to 7, with 1 indicating “very unlikely” and 7 indicating “very likely”. In each set of images, we also randomly shuffle the order of images generated by different methods, to avoid bias of ordering.
Figure 21 shows an example of such a set of images in the questionnaire. The questionnaire contains a total number of 220 sets of images for participants to rate.
Figure 21: Example of the questionnaire for human evaluation
Appendix E Details of Evaluation Setup
@@
For each illegal class and legal class in FF25 and the artwork dataset, we generally select 100 images in each class for mask learning, but if the number of images in the class is smaller than 150, we select half of the images for mask learning. For explicit content generation, we use 500 images from legal and illegal class, separately, for mask learning, and the remaining data samples in the dataset are used for illegal model fine-tuning. Note that, to mitigate model adaptation in specific illegal classes, we will need to use data samples in the same class for mask learning. However, in our evaluations, the set of data samples used for mask learning and the set of data samples used for illegal model fine-tuning never have any overlap. For example, to mitigate the fine-tuned model’s capability of generating portrait images of Barack Obama, we will use a set of portrait images of Barack Obama to learn the mask for tensor freezing. Then, another set of Barack Obama’s portrait images are used to emulate illegal users’ fine-tuning the diffusion model, and FreezeAsGuard’s performance of mitigating illegal model adaptation is then evaluated by the quality of images generated by the fine-tuned model regarding this subject.
@@
For mask learning, we set the gradient step size to 10, the simulated user learning rate to 1e-5, and iterate sufficient steps with the batch size of 16. The temperature for the mask’s continuous form is set to 0.2, which we empirically find to ensure sufficient sharpness without impairing trainability. When fine-tuning the diffusion model as an illegal user, we adopt a learning rate of 1e-5 and the batch size of 4 with Adam [36] optimizer. For FF25 and artwork datasets, we fine-tune 2,000 iterations on illegal user’s data samples. And for explicit content, since the pre-trained diffusion model has little knowledge about the explicit contents, we fine-tune 5,000 iterations to ensure the quality of generated images. Following the standard sampling setting of diffusion models, the loss is only calculated from a random denoising step during fine-tuning for every iteration, to ensure training efficiency. For image generation, we adopt the PNDMScheduler [33] and proceed with 50 denoising steps to ensure sufficient image quality.
@@
Appendix F More Qualitative Examples of Images Generated by the Fine-tuned Model
F.1 Forgery of Public Figures’ Portraits
@@
We provided more image examples in Figure 22, to show how FreezeAsGuard can effectively mitigate forgery of different public figures’ portraits. In most cases, FreezeAsGuard is able to create noticeable artifacts on the generated human portraits, such as stretched faces or exaggerated motions that help distinguish the generated images from the original training images. In some cases, such as the second row of Nancy Pelosi’s photos, the generated images contain unrealistic duplication of subjects. Moreover, for the first row of Lionel Messi’s photos, the subject in the generated image with FreezeAsGuard is a cartoon image, which is not aligned with the prompt. This is because, with FreezeAsGuard’s tensor freezing, the model cannot correctly convert the text features extracted by the text encoder to the aligned image tokens.
Figure 22: Examples of generated images after applying FreezeAsGuard-30% to Stable Diffusion v1.5 on illegal classes, where each prompt adopts the same seed for generation
F.2 Duplication of Copyrighted Artworks
Similarly, as more image examples in Figure 23 have shown, in most cases, images generated with baseline methods can exactly replicate the artistic style of the original training image. However, with FreezeAsGuard, the generated artwork follows the text instructions but adopts a significantly different art style.
Figure 23: Examples of generated images after applying FreezeAsGuard-70% to Stable Diffusion v2.1 on illegal classes, where each prompt adopts the same seed for generation
F.3 Generation of Explicit Contents
As shown in Figure 24, the generated images with FreezeAsGuard can effectively avoid explicit contents from being shown in different ways. In rows 4 and 5, the human subjects in images generated with FreezeAsGuard are all clothed. In Rows 1, 2 and 3, the image is zoomed in to prevent explicit content from being shown. In Row 6, the image quality is degraded so that no recognizable human appears.
Figure 24: Examples of generated images after applying FreezeAsGuard-70% to Stable Diffusion v1.4 on illegal classes, where each prompt adopts the same seed for generation
Appendix G Ethical Issues of Using the Public Portrait Images and Artwork Images
In this section, we affirm that the use of our self-collected public portrait images and artwork image dataset does not raise ethical issues.
G.1 Image Source
For the FF-25 dataset, we use the Google images search API to crawl the images from the Web. Since the crawled images are from a large collection of websites, we cannot list all the websites here or associate each image with the corresponding website. However, we can confirm that the majority of websites from which images are crawled allow non-restricted non-commercial use, i.e., the CC NC or CC BY-NC license. Some examples of these websites are listed as follows:
  • Wikipedia.org
  • whitehouse.gov
  • ifeng.com
  • theconversation.com
  • house.gov
  • cartercenter.org
  • newstatesman.com
  • esportsobserver.com
  • slate.fr
  • letemps.ch
For the artwork image dataset, we use artist’s posted images on their public Instagram accounts. The following keywords can be used to search these public Instagram accounts:
  • Beeple_crap
  • Saonserey
  • Kylelambertartist
  • Davidsossella
  • Thebutcherbilly
G.2 Image Usage
Our collection and use of these images are strictly limited to non-commercial research use, and these images will only be released to a small group of professional audience (i.e., CVPR reviewers) instead of the wide public. Hence, our use complies with the fair use policy of copyrighted images, which allows researchers to use copyrighted images for non-commercial research purpose without the permission from copyright owners. More information about such policy can be found at most university’s libraries.
G.3 Use Policy in the Research Community
We noticed that such fair use policy mentioned before has been widely applied in the research community to allow usage of copyrighted images of public figures’ portraits and artworks for research purposes. For example, many datasets of celebrities’ portraits such as CelebA [42], PubFig [37] and MillionCelebs [63]) and artwork such as Wikiart [53] and LION [49] are publicly available online. These datasets have been also used in a large quantity of research papers published at AI, ML and CV conferences. For examples: [35, 18] used the CelebA dataset, [32, 34] used the PubFig dataset and [58] use the WikiArt dataset.